Phase transitions are similar to chemical reactions in that they each have an associated entropy change. A phase transition involves the breaking or forming of intermolecular attractive forces.

A change in entropy accompanies not only a variation in the temperature, pressure, or volume of a system but also a physical transformation such as the fusion, vaporization, or transition from one crystalline state into another.

In the equilibrium between two phases, the exchange of heat can be carried out reversibly, and the quantity of heat is equal to the enthalpy change for the transition, ΔHtr. The associated change in entropy, ΔStr, is given by the equation

ΔStr=ΔHtrTtr

where Ttr is the transition temperature.

It has been observed that most organic liquids exhibit almost the same value for the entropy of vaporization. This observation is described by Trouton's rule and is useful for making predictions about liquids.

Predict which of the following compounds follow Trouton's rule.

Check all that apply.Part B Enthalpy of Boiling point Liquid moly 17.5 383.6 423.8 349.8 36.24 33.18 37.19 2982 38.58 18.67 loluene carton tetrachloride ethanol hydrogen sulfride 213.45 Prodict which of the following compound& follow Troutons ruo. Check all that apply. View Available Hints) ethanol cluene hydrogen suifide Submit Previous Answors
Answer #1

The ΔSvap is calculated according to the Trouton rule of all the compounds and the results are compared:

1) O-xylene: ΔS = 36240 / 417.5 = 86.8 J / mol * K

2) Toluene: ΔS = 33180 / 383.6 = 86.5 J / mol * K

3) Nonano: ΔS = 37180 / 423.8 = 87.73 J / mol * K

4) CCl4: ΔS = 29820 / 349.8 = 85.25 J / mol * K

5) Ethanol: ΔS = 38560 / 351.3 = 109.76 J / mol * K

6) Hydrogen Sulfide: ΔS = 18670 / 213.45 = 87.47 J / mol * K

According to Trouton's rule, compounds in which their ΔSvap are between 85 and 88 J / mol * K, are under it.

O-xylene, Toluene, Nonano, CCl4 and Hydrogen Sulfide are under Trouton's rule.